Sheridan Furniture Company started construction of a combination office and warehouse building for its own use at an estimated cost of $4,972,800 on January 1, 2017. Sheridan expected to complete the building by December 31, 2017. Sheridan has the following debt obligations outstanding during the construction period.

Construction loan-10% interest, payable semiannually, issued December 31, 2016 $2,019,500
Short-term loan-8% interest, payable monthly, and principal payable at maturity on May 30, 2018 1,605,600
Long-term loan-9% interest, payable on January 1 of each year. Principal payable on January 1, 2021 1,001,000

Compute the depreciation expense for the year ended December 31, 2018. Sheridan elected to depreciate the building on a straight-line basis and determined that the asset has a useful life of 30 years and a salvage value of $302,700. (Round answer to 0 decimal places, e.g. 5,275.)

Depreciation Expense

$
Answer #1
CAPITALIZATION OF INTEREST :
A Accumulated expenditure in beginning 2017 $0
B Accumulated expenditure at end 2017 $4,972,800
C=(A+B)/2 Average accumulated expenditure $2,486,400
D Construction Loan $2,019,500
E=C-D Other Loan $466,900
F Interest on construction loan=10%*2019500 $201,950
OTHER LOANS
Loan amount Percent Annual Interest
Short Term loan $1,605,600 8% $128,448
Long termloan $1,001,000 9% $90,090
TOTAL $2,606,600 8.38% $218,538
Weighted average interest rate=218538/2606600 8.38%
Interest Cost tobe capitalized<
Borrowing amount Interest rate Interest cost
Construction Loan $2,019,500 10% $201,950
Other borrowing $466,900 8.38% $39,145
Total $241,095
Asset to be depreciated :
Accumulated expenditure at end 2017 $4,972,800
Capitalized Interest $241,095
G Total capitalized asset $5,213,895
H Salvage Value $302,700
I Useful Life in years 30
AnnualDepreciation =(Capitalized asset-Salvage Value)/(Usefullife)
J=(G-H)/I Depreciation expense $163,707
l